OpenStudy (anonymous):

a water tunnel has a circular cross section where the diameter diminishes from 3.6 to 1.2 if the velocity of water is 3.0 m/s in the larger part of the tunnel what is the velocity of water in the smaller part of the tunnel

OpenStudy (anonymous):

Bernoulli's equation will give you the answer.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Bernoulli's_principle

OpenStudy (anonymous):

27m/s ? I believe this is it

OpenStudy (anonymous):

Hmmm, I'm not sure what level of physics you are at, but I would say 12 m/s

OpenStudy (anonymous):

AS the flow rate remain constant so the answer is 27 m/sec

OpenStudy (anonymous):

Using the equation of continuity of flow, \(A_1V_1=A_2V_2\) \(\pi (r_1)^2 V_1=\pi (r_2)^2 V_2\) \((r_1)^2 V_1=(r_2)^2 V_2\) plug it all in, \((3.6)^2 (3)=(1.2)^2 V_2\) \(V_2=27ms^{-1}\)